OpenStudy (anonymous):

Rewrite the expression without using grouping symbols. –6(8 + 12) A. (–6) • 8 + 12 B. (–6) • 8 + (–6) • 12 C. –68 + 612 D. (–6) + 8 • (–6) + 12

OpenStudy (mathstudent55):

Distributive property: a(b + c) = ab + bc -6(8 + 12) = -6(8) + (-6)(12)

OpenStudy (anonymous):

huh?

OpenStudy (mathstudent55):

What don't you understand?

OpenStudy (mathstudent55):



OpenStudy (anonymous):

sooooo

OpenStudy (anonymous):

b

OpenStudy (thebadyboy):

correct